# On Strong Lefschetz Property of 0-dimensional complete intersections

### v2

@@ Abstract.
We prove that a homogeneous 0-dimensional complete intersection satisfies the Strong Lefschetz Property (SLP) in degree 1 if and only if its associated form has nonzero Hessian.
@@ 1. Statements of main results
where the action of $S$ on $R$ is obtained by identifying $x_{i}$ with $\partial/\partial y_{i}$. Moreover, for any finitely generated graded Artinian Gorenstein $\mathbb{C}$-algebra $S/J$ with socle degree $T$, there exists a homogeneous polynomial $F\in R_{T}$, unique up to scaling, such that $J=I_{F}$; see [6, Lemma 2.12]. Any such $F$ is called a Macaulay inverse system for $S/J$. It is known that the associated form $A_{\bf f}(y_{1},\ldots,y_{n})\in R_{T}$ is a Macaulay inverse system for the Artinian algebra $M({\bf f})$, see [1, Proposition 3.2].
@@
As in [4, Remark 1.2], let $f\in S_{d}$ be a reduced homogeneous polynomial of degree $d\geq 3$. By Euler’s formula, the singular subscheme of the reduced hypersurface $X=V(f)\subseteq\mathbb{P}^{n-1}$ is defined by the gradient ideal
@@
We would like to thank A. Dimca for useful remarks. The author is supported by NSFC-12301100.
@@ 2. Proof
Recall that $J({\bf f})=\mathrm{Ann}(\mathrm{A}_{\bf f})$. By [6, Lemma 2.15], we have $J({\bf f})_{T-1}=\mathrm{Ann}(\mathrm{A}_{\bf f})_{T}:S_{1}$, where

### v3

@@ Abstract.
We prove that a homogeneous 0-dimensional complete intersection satisfies the Strong Lefschetz Property (SLP) in degree 1 if and only if its associated form has nonzero Hessian. The result is essentially known in the literature, but our proof is different compared with the previous ones.
@@ 1. Statements of main results
where the action of $S$ on $R$ is obtained by identifying $x_{i}$ with $\partial/\partial y_{i}$. Moreover, for any finitely generated graded Artinian Gorenstein $\mathbb{C}$-algebra $S/J$ with socle degree $T$, there exists a homogeneous polynomial $F\in R_{T}$, unique up to scaling, such that $J=I_{F}$; see [7, Lemma 2.12]. Any such $F$ is called a Macaulay inverse system for $S/J$. It is known that the associated form $A_{\bf f}(y_{1},\ldots,y_{n})\in R_{T}$ is a Macaulay inverse system for the Artinian algebra $M({\bf f})$, see [1, Proposition 3.2].
@@
As in [5, Remark 1.2], let $f\in S_{d}$ be a reduced homogeneous polynomial of degree $d\geq 3$. By Euler’s formula, the singular subscheme of the reduced hypersurface $X=V(f)\subseteq\mathbb{P}^{n-1}$ is defined by the gradient ideal
@@
We were made aware that our result has appeared in prior works (see e.g. [8, Theorem 3.1],[4, Corollary 2.14 (2)]). Thus, this note gives a self-contained proof of a known result via a different approach.
@@ 2. Proof
Recall that $J({\bf f})=\mathrm{Ann}(\mathrm{A}_{\bf f})$. By [7, Lemma 2.15], we have $J({\bf f})_{T-1}=\mathrm{Ann}(\mathrm{A}_{\bf f})_{T}:S_{1}$, where